# Bayesian spatiotemporal modelling of political violence and conflict events using discrete-time Hawkes processes

### v1

@@ 5.1 Summary
In this analysis we presented a unified framework using spatiotemporal DTHPs to analyse both the short and long term trends of conflict risk. Through the Bayesian framework, we can gain a more precise understanding of the uncertainty inherent in our estimates of conflict risk, compared to other approaches based on machine learning techniques and optimisation methods such as maximum likelihood estimation. To our knowledge this is the first analysis of its kind applied to the ACLED database, and the first to compare the various conflict types between each other and also among neighbouring countries. We also demonstrated from the practitioners’ perspective the advantages of our approach and examples of how to interact with our model to gain valuable insights.
@@
In response to our first research question, we find that spatiotemporal DTHPs are indeed a useful tool for policy makers to model and monitor conflict events. Our model was fitted to each of the individual countries within South Asia, and broken down further by the conflict type. The resulting models performed well and reacted quickly to recent events to capture and understand the self-exciting nature of conflict in the short term. A less computationally intensive approach based on maximum likelihood estimation was also presented, and we determined that while this approach gives reasonable point estimates, the Bayesian model is preferred due to more reliable uncertainty quantification. We also performed out-of-sample predictive checks, which showed that our model was generally capable of producing a reasonable range of estimates for the predicted number of events over a 3 month period, with the exception of some extreme observations.
@@
There are several key learnings from using spatiotemporal DTHPs for modelling conflict events that address our second question. While other modelling approaches can also provide a measure of conflict risk, interpretation of the parameters in our model offers unique insights. The level of independent events and those that are caused by past events can be directly quantified through the baseline and magnitude parameters respectively. This can help to inform preventative measures, for example where the baseline rate of events is high, there should be a focus on preventing new events from occurring. Whereas, if the magnitude parameter is high, there should be focus on monitoring ongoing events. We also quantified the effect that time and distance have on the risk of future conflict events by explicitly parameterising the temporal and spatial patterns of the process through their respective triggering kernels. We then used these kernels to compare the self-exciting characteristics of the various conflict types and countries to further deepen our understanding of these processes.
@@
To address our final research question, we provided specific examples of how to use and adopt our models in practice and benchmark against current practices. We demonstrated how an early warning tool, used to monitor longer-term trends in the data and thus inform anticipatory actions, can be developed based on our statistical model. Specifically, this tool identifies when the observed level of events exceeds a threshold percentile determined by the posterior distribution, and we showed that our model is more stable and robust than the alternative based on historical averages. Visualisations for the current level of conflict risk estimated from our model were also presented. Our modelling was performed at a finer spatial resolution compared to similar analyses, allowing for more accurate estimates of conflict risk, even when aggregating into coarser spatial regions. Lastly, our statistical modelling approach allows us to make short-term predictions, and calculate the corresponding uncertainty in these estimates.
@@ 5.2 Limitations and future work
Given sufficient computational resources or a more computationally efficient inference method, a global, multivariate model for conflict events could be estimated to characterise the interactions between countries and regions. Alternatively, if interactions were less important, and instead similarities between regions were of interest, univariate studies could be conducted in parallel and the estimated parameters grouped via a clustering algorithm to determine countries or regions with similar risk profiles. Furthermore, a multivariate model between conflict types could be fitted, to capture and quantify dependencies between each of the conflict types. The level of risk identified in our analysis could also be compared to other trends, for example economic or demographic time series to determine any interactions.
The estimation of the spatial kernel enables the characterisation of spatial risk profiles between countries and conflict types. However, its inclusion does not greatly improve predictive accuracy as shown in the supplementary materials. While prediction is not a key focus of this article, further work could be undertaken to explore alternative modelling approaches to improve predictive performance.
@@
One phenomenon observed throughout this analysis was that the proposed model struggles to capture more extreme event counts. The mean intensity, represented by $\lambda(t)$, will rarely reach these peaks since they are in the tails of the posterior distribution. However, some modifications could be made to the model to better represent this range in values. For example, one could incorporate an over-dispersion term by adopting a negative binomial likelihood as opposed to the Poisson likelihood. Another approach could use a hurdle model (Porter and White, 2012), that models a binary response for whether there will be a non-zero number of events, and the magnitude of events given a non-zero count. Further extensions could be considered to accommodate excess counts under this framework, such as a hidden Markov model or a simple mixture model.
@@
The level of discretisation, both spatial and temporal, is also an important consideration in the modelling. The inference is more stable given a coarser discretisation as it avoids the issue of zero-inflated event counts. A coarser aggregation level also reduces the computational expense of the inference. However, there is an approximation error introduced in the discrete-time model that assumes events that occur in the same interval are independent, and thus cannot have been produced by one another. Therefore, if the discretisation is too coarse, there is the potential to artificially inflate the baseline rate. The discretisation error is explored for temporal Hawkes processes in Kirchner (2016). Further work could be undertaken to explore the impact of this approximation error, and to determine the optimal discretisation level for a given dataset. The same arguments also apply to the spatial aggregation, however in this case study we believe the administrative boundaries used are sensible given the implicit modelling assumption of homogeneity between all events that occur in the same region. Siviero et al. (2024) also investigate this discretisation error for spatiotemporal Hawkes processes.

### v2

@@ 5.1 Summary
In this analysis we presented a unified framework using spatiotemporal DTHPs to analyse both the short and long term trends of conflict risk. Through the Bayesian framework, we can gain a more precise understanding of the uncertainty inherent in our estimates of conflict risk, compared to other approaches based on machine learning techniques and optimisation methods such as maximum likelihood estimation. To our knowledge this is the first analysis of its kind applied to the ACLED database, and the first to characterise and compare the spatial and temporal dynamics of the various conflict types between each other and also among neighbouring countries. We also demonstrated from the practitioners’ perspective the advantages of our approach and examples of how to interact with our model to gain valuable insights.
@@
In response to our first research question, we find that spatiotemporal DTHPs are indeed a useful tool for policy makers to model and monitor conflict events. Our model was fitted to each of the individual countries within South Asia, and broken down further by the conflict type. The resulting models performed well allowing us to capture and understand the self-exciting nature of conflict in the short term. We also performed out-of-sample predictive checks, which showed that our model was generally capable of producing a reasonable range of estimates for the predicted number of events over a 3 month period, with the exception of some extreme cases.
@@
There are several key learnings from using spatiotemporal DTHPs for modelling conflict events that address our second question. While other modelling approaches can also provide a measure of conflict risk, interpretation of the parameters in our model offers unique insights. The level of independent events and those that are caused by past events can be quantified through the baseline and magnitude parameters respectively. This can help to inform preventative measures, for example where the baseline rate of events is high, there should be a focus on preventing new events from occurring. Whereas, if the magnitude parameter is high, there should be focus on monitoring ongoing events. We also fitted several types of spatiotemporal DTHP and investigated whether a spatially varying baseline rate alone was sufficient to the characterise spatial dynamics in the data, or if spatial excitation was a factor in driving conflict events. In doing so, we quantified the effect that time and distance have on the risk of future conflict events by explicitly parameterising the temporal and spatial patterns of the process through their respective triggering kernels. We then used these kernels to compare the self-exciting characteristics of the various conflict types and countries to further deepen our understanding of these processes.
@@
To address our final research question, we provided specific examples of how to use and adopt our models in practice and benchmark against current practices. We demonstrated how an early warning tool, used to monitor longer-term trends in the data and thus inform anticipatory actions, can be developed based on our statistical model. Specifically, this tool identifies when the observed level of events exceeds a threshold percentile determined by the posterior distribution, and we showed that our model is more stable and robust than the alternative based on historical averages. Visualisations for the current level of conflict risk estimated from our model were also presented. Our modelling can be performed at a finer spatial resolution compared to similar analyses, allowing for more accurate estimates of conflict risk, even when aggregating into coarser spatial regions. Lastly, our statistical modelling approach allows us to make short-term predictions, and calculate the corresponding uncertainty in these estimates.
@@ 5.2 Limitations and future work
Given sufficient computational resources or a more computationally efficient inference method, a global, multivariate model (Hawkes, 1971; Li and Cui, 2024) for conflict events could be estimated to characterise the interactions between countries and regions. Alternatively, if interactions were less important, and instead similarities between regions were of interest, univariate studies could be conducted in parallel and the estimated parameters grouped via a clustering algorithm to determine countries or regions with similar risk profiles. Furthermore, one could extend the spatiotemporal DTHP to a marked process whereby the type of conflict event is a mark which also influences the excitation dynamics (Ogata, 1988; Alaimo Di Loro et al., 2025). A more complete option which allows for different triggering kernels for each pair of conflict types is a multivariate model between conflict types to capture and quantify dependencies between each of the conflict types. The level of risk identified in our analysis could also be compared to other trends, for example economic or demographic time series to determine any interactions.
In the present work, we incorporated a spatially varying baseline rate to better capture spatial heterogeneity in event dynamics. This extension allows for more realistic modelling of the underlying spatial structure while maintaining computational feasibility. Several further developments could be explored in future work. First, it may be possible to extend branching structure–based approaches (Alaimo Di Loro et al., 2025; D’Angelo et al., 2024; Fox et al., 2016; Zhuang and Mateu, 2018) to the discrete-time Hawkes framework, though this remains technically challenging due to the absence of a clear branching interpretation in discrete time. Additionally, the baseline rate could be refined through regression-based adjustments using covariates such as population density or other socio-demographic factors of interest. Alternative modelling strategies include adopting a multivariate (regional) formulation as discussed above or employing Gaussian Process-based methods, such as the Log Cox Gaussian Process (Miscouridou et al., 2023), for encoding spatial variation.
In this study, Euclidean centroid-to-centroid distances based on longitude and latitude were used as a measure of spatial separation. Given that the study countries lie relatively close to the equator and the administrative units are generally small, this approximation is reasonable for our purposes. However, we acknowledge that this approach becomes less appropriate in settings involving larger spatial domains or regions at latitudes further from the equator. In such contexts, projections to suitable coordinate reference systems, for example UTM, or the use of other distances, such as Haversine distances, would provide more accurate distance representations, although care is needed when countries span multiple UTM zones. Furthermore, centroid-based distances may oversimplify interactions between neighbouring areas, and alternatives such as adjacency-based structures or shared border lengths could offer more meaningful representations of spatial connectivity. Finally, extensions of continuous-space methodologies, such as those used in the discrete approximations of LGCPs (Li et al., 2012; Johnson et al., 2019), may offer a route toward more flexible spatial interaction modelling within the Hawkes process framework.
@@
One phenomenon observed throughout this analysis was that the proposed model does not always capture more extreme event counts. However, if it were reasonable to assume that these excess counts have some inherent structure, some modifications could be made to the model to better represent this range in values. For example, one could incorporate an over-dispersion term by adopting a negative binomial likelihood as opposed to the Poisson likelihood. Another approach could use a hurdle model (Porter and White, 2012), that models a binary response for whether there will be a non-zero number of events, and the magnitude of events given a non-zero count. Further extensions could be considered to accommodate excess counts under this framework, such as a hidden Markov model or a simple mixture model to capture varying levels of intensity.
Due to the model formulation there is a systematic delay in the modelled risk estimates whereby events that occur in the current time period can only influence the intensity of the process in the subsequent time period. To offset this delay, a potential future direction could consider covariates derived from topic modelling of newspaper articles (Mueller and Rauh, 2022) or other exogenous, real-time covariates. This information could mitigate this delayed effect by allowing the model to respond to external drivers that correlate with imminent changes in event frequency.
@@
The level of discretisation, both spatial and temporal, is also an important consideration in the modelling. The inference is more stable given a coarser discretisation as it avoids complications arising from excessive zero-inflated event counts, and it also reduces the computational expense of the inference. However, there is an approximation error introduced in the discrete-time model that assumes events that occur in the same interval are independent, and thus cannot have been produced by one another. Therefore, if the discretisation is too coarse, there is the potential to artificially inflate the baseline rate. The discretisation error is explored for temporal Hawkes processes in Kirchner (2016a). Further work could be undertaken to explore the impact of this approximation error, and to determine the optimal discretisation level for a given dataset. The same arguments also apply to the spatial aggregation, however in this case study we believe the administrative boundaries used are sensible given the implicit modelling assumption of homogeneity between all events that occur in the same region. Siviero et al. (2024) also investigate this discretisation error for spatiotemporal Hawkes processes.
In this work we adopted general forms of the modelling assumptions across all scenarios, such as the choice of triggering kernels and aggregation levels, to assess the suitability of spatiotemporal DTHPs for analysing conflict activity. This analysis can then inform specific scenarios which may require further refinement of the model assumptions; for example, a different temporal aggregation may be more appropriate depending on the research question. We expect that it will be useful for users to be able to tailor these assumptions to their specific scenarios. As such, a more automated and streamlined workflow for determining the key modelling decisions, such as the form of $\lambda$, maximum time lag, choice of kernels, temporal and spatial aggregation level and choice of likelihood distribution, should be developed. It would also be helpful for assessing goodness-of-fit to extend the existing, formal residual analyses available for continuous-time Hawkes processes (Ogata, 1988; Daley and Vere-Jones, 2006) to the discrete-time setting.